# Algorithmic Information Theory for Graph Edge Grouping and Substructure Analysis

### v2

@@ Introduction
We can assign a probability value to a string $s$ by running the Busy Beaver function for $t$ states with $k$ symbols (0,1 for binary sequences). Instead of counting the number of 11s produced by a Turing machine, we count the number of Turing machines that produce $s$ over the number of all Turing machines for $t$ states and $k$ symbols that halt.
@@
We now fix the values of $t$ and $k$ to have a specific distribution table based on $D(t,k)$, which is represented as CTMCTM based in Equation 5. Let $l$ be an integer such that $l\leq s$ and is the length of each substring $s^{i}$ when $s$ is decomposed (with a possible remainder of $y=s\mod l$ characters). After decomposition, there is a probability that some $s^{i}$ substrings are the same, thus $n_{i}$ is the multiplicity of each $s^{i}$ found. If $p_{i}$ is the smallest program that produces substring $s^{i}$, then $\log(n_{i})$ is the number of bits needed to encode $p_{i}$.
@@
Each element in $\lambda$ is a permutation of $V(G)$ such that when $V(G)$ is permuted by any element in this set, it is automorphic to any other permutation in the same set. This ultimately groups each permutation $\gamma\in S_{G}$ to a specific automorphic subset. Since a graph can have a number of automorphic subsets, we define the complete automorphic set of a graph as:
@@
$\mathfrak{C}$ (Algorithm 2) uses a set of permutations $\Gamma$ to permute the vertices of a graph $G$ and perform edge perturbation. Initially, a set of tuples II is created where the average algorithmic information contribution of each edge $(u,v)\in E_{G}$ will be housed and can be retrieved by using $(u,v)$ as a key in II which is denoted by $I[(u,v)]$. Using a permutation $\gamma\in\Cup(G)$, we permute $G$ and retrieve its adjacency matrix $X$ that is denoted by $Adj(E_{G^{\gamma}})$. The row and column of the matrix are arranged according to the ordered labelling using $V_{G^{\gamma}}$, where each $v\in V_{G^{\gamma}}$ is the $v$th row/column in the matrix $X$.
For each permuted edge $(w,x)=(\gamma(u),\gamma(v))$ where $u,v\in V_{G}$, we compute its information contribution value $\text{bdm}_{(w,x)}$ by removing $(w,x)$ from $G^{\gamma}$ which is denoted by $E_{G^{\gamma}}\;\backslash\;(w,x)$. After removal, the difference in $BDM_{2D}$ between the original matrix $X$ and the matrix $Y$ (where the edge is removed) is added to the existing value stored in II. To properly add the information contribution for the edges when permuted by $\gamma$, we revert $(w,x)$ to its original label in $E_{G}$ by the inverse function $\gamma^{-}$ since calculating the information contribution of each edge is done in each permutation $\gamma\in\Cup(G)$. Finally, each information contribution value of an edge in II is averaged by the number of $\gamma\in\Cup(G)$ since $|\Cup(G)|=|\Lambda(G)|$.
@@ Results
where $O(1)$ accounts for the constant program size to connect two vertices from each sub-graph by an edge whose program size is independent of $G$. Let $G^{*}$ be the shortest program that produces $G$ such that $K(G)=|G^{*}|$. If $G$ has two subgraphs, then $|G^{*}|\leq|G^{*}_{1}|+|G^{*}_{2}|+O(1)$. When computing for $K(G)$, if edges $e_{i},e_{j}\in E_{G_{1}}$, then the bits needed to encode $e_{i}$ and $e_{j}$ into $G$ are included in $G_{1}^{*}$. There are three main modes of inquiry that can be asked to determine the effectivity of edge perturbation using BDMBDM:
@@
To further determine the effectiveness of edge perturbation, we also analyzed how similar the information contribution of an edge to other edges. Since we have computed the average information contribution for every edge of a graph $G$ using $\mathfrak{C}$, we can then compare if edges per sub-graph ($G_{1}$ & $G_{2}$) have grouped correctly using the sorted set of tuples II (recall that II is arranged from highest to lowest in average information contribution value). When $\mathfrak{C}$ outputs II, it shows the likelihood of each edge to be grouped with other edges that were produced by the same underlying sub-program. We categorize each graph into three different grouping schemes with respect to the linear arrangement of the average information contribution values of $E_{G}$:
@@
    Complete - all the edges in II have grouped $G_{1}$ and $G_{2}$’s edges respectively.
@@
Averaging of information contribution values has been done because different vertex permutations result in different $K(G)$ values. Since $\Lambda(G)$ groups isomorphisms of different vertex permutations that are automorphic per grouping, $\Cup(G)$ is the compressed version of $\Lambda(G)$ and ultimately of $S_{G}$. Pairing this approach of getting unique permutations and averaging of each edge’s information contribution in a graph allows us to have a fair assessment of BDMBDM by probabilistic and combinatorial means. We computed the average information contribution values with respect to the parameter $\Gamma$ in $\mathfrak{C}(G,\Gamma)$ to determine if $\Cup(G)$ is already effective without using $S_{G}$ since $\Cup(G)$ is algorithmically quicker to execute. Tables 2 and 3 show a comparison of $\Cup(G)$ and $S_{G}$. We can see that the symmetric group is more effective compared to using automorphic subset sampling. Although using $\Cup(G)$ has partially grouped edges of some graphs, $S_{G}$ correctly grouped edges for more graphs.
@@
The connecting edge does not have $\max_{\text{info}}$ using this permutation set. Let $e_{n}$ be the connecting edge which is placed in the $n$th location in $I_{k}=\{e_{n+1},e_{n},e_{n-1},\ldots,e_{1}\}$, where $I_{k}$ is the set of keys for the set of tuples II and $I[e_{n}]$ is the average information contribution value of the connecting edge. If the connecting edge is $\max_{\text{info}}$, then $n=|I|$ (the first value in the linear arrangement).
As seen in Table 4, not only did $\mathfrak{C}$ identify that the connecting edge was $\max_{\text{info}}$} for most graphs using both $\Gamma=\Cup(G)$ and $\Gamma=S_{G}$, but it also managed to identify the distance of the connecting edge to the next edge in the linear arrangement of average information contribution was more than the cutoff for half of the graphs. Out of the 30 graphs, 16 of them satisfied $e_{n}-e_{n-1}>\log_{2}(2)$. Interestingly, it shows that performing averaging of information contribution with either $\Cup(G)$ or $S_{G}$ was effective when graphs contain a sub-graph that has a high edge count. Among these 16 graphs, either a Watts-Strogatz or a Complete graph was present as a sub-graph for each graph. This indicates that BDMBDM is effective in detecting tightly connected substructures within a graph. Further proof is that when a graph only contains these two subgraphs, the connecting edge’s average information contribution is high relative to the other graphs examined.
Conversely, BDMBDM did not satisfy $e_{n}-e_{n-1}>\log_{2}(2)$ for 14 out of the 30 graphs (colored red in Table 4) when $\Gamma=\Cup(G)$. 6 of the 9 graphs have a random sub-graph that is included per graph. When both subgraphs are random graphs, then adding a connecting edge will not be distinguishable from the other edges in the graph because the entire graph would be considered one combined random graph. There was not much change in $e_{n}-e_{n-1}$ when using $S_{G}$ over $\Cup(G)$ except for graphs that had a cycle sub-graph. This indicates that averaging of information contribution using $S_{G}$ provides a more fine-grained inspection of substructures. Appending to the findings in the Edge Grouping section of the results, most of the graphs that were classified as complete or partial edge grouping consist of mostly graphs with regular and small-world subgraphs, namely Cycle, Complete and Watts-Strogatz graphs.
@@ Discussion
There are 2 limitations that prohibit us from further experimentation of larger graphs. First is that $\mathfrak{C}$ grows by $n!$, where $n$ is the number of nodes of the graph. To fully test BDMBDM on the basis of combinatorial and probabilistic means, much bigger graphs should be tested using $\mathfrak{C}$. Although $\mathfrak{C}$ shows promise of identifying substructures in a graph and was able to:
@@
it must be tested in real world networks whose number of nodes is greater than 12 to have a more accurate testing of BDMBDM. The second limitation is in the sub-matrix size that CTMCTM (within BDMBDM) can slice an adjacency matrix of a graph. Currently, BDMBDM can handle $3\times 3$ or $4\times 4$ sub matrices and uses the periodic partitioning when $n$ is not divisible by $l$. Ideally, since BDMBDM is an approximated upper-bound of algorithmic complexity, we would like for CTMCTM to handle $l>4$ for larger graphs since there is an error overhead whenever $n$ is not divisible by $l$ and thus the remaining rows and columns in the adjacency matrix are appended with the previous rows and columns using a partitioning technique. But this is a monumental task since computing for larger matrices in CTMCTM is equivalent to running the Busy Beaver function.
@@ Conclusion
A combination of permuting a graph then averaging the information contribution per edge perturbed was done to 30 synthetic graphs to determine the efficacy of BDMBDM as a method in revealing substructures within complex networks. The sets of permutations that the algorithm $\mathfrak{C}$ (which uses BDMBDM) used was the automorphic subsets $\Cup(G)$ of a graph and its symmetric group $S_{G}$. We used these two permutation sets to compare a compressed representation of a graph’s structure (automorphic subsets) from its entire permutation set (symmetric group). Out of the 30 graphs, 29 of them have been successfully identified as having the connecting edge as the edge with the highest average information contribution. Within these 29 graphs, 16 were also identified of having the connecting edge as the edge with a distance of more than $\log_{2}(2)$ from other edges within their respective graphs. To highlight, edge perturbation was proven to be effective for graphs that have tightly connected subgraphs, namely Watts-Strogatz graphs and Complete graphs. The connecting edge was clearly differentiated in terms of its average information contribution from the edges of the subgraphs. Averaging of information contribution using $\Cup(G)$ and $S_{G}$ was also effective in completely or partially grouping edges together. Comparing the two permutation sets, $S_{G}$ performed better since more graphs’ edge grouping has classified as a complete edge grouping.
@@ Acknowledgments
I would like to acknowledge my thesis adviser Briane Samson for guiding and motivating me when writing this paper. I would also like to thank Hector Zenil for answering my inquiries regarding their initial implementation of causal deconvolution along with Allan Zea, who implemented the R package of causal deconvolution. Lastly, I would like to show my gratitude to the maintainer of the Python package that implements BDMBDM, Szymon Talaga, for answering questions regarding implementing edge perturbation to the package.

### v3

@@ Introduction
We can assign a probability value to a string $s$ by running the Busy Beaver function for $t$ states with $k$ symbols (0,1 for binary sequences). Instead of counting the number of $1$s produced by a Turing machine, we count the number of Turing machines that produce $s$ over the number of all Turing machines for $t$ states and $k$ symbols that halt.
@@
We now fix the values of $t$ and $k$ to have a specific distribution table based on $D(t,k)$, which is represented as $CTM$ based in Equation 5. Let $l$ be an integer such that $l\leq s$ and is the length of each substring $s^{i}$ when $s$ is decomposed (with a possible remainder of $y=s\mod l$ characters). After decomposition, there is a probability that some $s^{i}$ substrings are the same, thus $n_{i}$ is the multiplicity of each $s^{i}$ found. If $p_{i}$ is the smallest program that produces substring $s^{i}$, then $\log(n_{i})$ is the number of bits needed to encode $p_{i}$.
@@
Each element in $\lambda$ is a permutation of $V_{G}$ such that when $V_{G}$ is permuted by any element in this set, it is automorphic to any other permutation in the same set. This ultimately groups each permutation $\gamma\in S_{G}$ to a specific automorphic subset. Since a graph can have a number of automorphic subsets, we define the complete automorphic set of a graph as:
@@
$\mathfrak{C}$ (Algorithm 2) uses a set of permutations $\Gamma$ to permute the vertices of a graph $G$ and perform edge perturbation. Initially, a set of tuples $I$ is created where the average algorithmic information contribution of each edge $(u,v)\in E_{G}$ will be housed and can be retrieved by using $(u,v)$ as a key in $I$ which is denoted by $I[(u,v)]$. Using a permutation $\gamma\in\Cup(G)$, we permute $G$ and retrieve its adjacency matrix $X$ that is denoted by $Adj(E_{G^{\gamma}})$. The row and column of the matrix are arranged according to the ordered labelling using $V_{G^{\gamma}}$, where each $v\in V_{G^{\gamma}}$ is the $v$th row/column in the matrix $X$.
For each permuted edge $(w,x)=(\gamma(u),\gamma(v))$ where $u,v\in V_{G}$, we compute its information contribution value $\text{bdm}_{(w,x)}$ by removing $(w,x)$ from $G^{\gamma}$ which is denoted by $E_{G^{\gamma}}\;\backslash\;(w,x)$. After removal, the difference in $BDM_{2D}$ between the original matrix $X$ and the matrix $Y$ (where the edge is removed) is added to the existing value stored in $I$. To properly add the information contribution for the edges when permuted by $\gamma$, we revert $(w,x)$ to its original label in $E_{G}$ by the inverse function $\gamma^{-}$ since calculating the information contribution of each edge is done in each permutation $\gamma\in\Cup(G)$. Finally, each information contribution value of an edge in $I$ is averaged by the number of $\gamma\in\Cup(G)$ since $|\Cup(G)|=|\Lambda(G)|$.
@@ Results
where $O(1)$ accounts for the constant program size to connect two vertices from each sub-graph by an edge whose program size is independent of $G$. Let $G^{*}$ be the shortest program that produces $G$ such that $K(G)=|G^{*}|$. If $G$ has two subgraphs, then $|G^{*}|\leq|G^{*}_{1}|+|G^{*}_{2}|+O(1)$. When computing for $K(G)$, if edges $e_{i},e_{j}\in E_{G_{1}}$, then the bits needed to encode $e_{i}$ and $e_{j}$ into $G$ are included in $G_{1}^{*}$. There are three main modes of inquiry that can be asked to determine the effectivity of edge perturbation using $BDM$:
@@
To further determine the effectiveness of edge perturbation, we also analyzed how similar the information contribution of an edge to other edges. Since we have computed the average information contribution for every edge of a graph $G$ using $\mathfrak{C}$, we can then compare if edges per sub-graph ($G_{1}$ & $G_{2}$) have grouped correctly using the sorted set of tuples $I$ (recall that $I$ is arranged from highest to lowest in average information contribution value). When $\mathfrak{C}$ outputs $I$, it shows the likelihood of each edge to be grouped with other edges that were produced by the same underlying sub-program. We categorize each graph into three different grouping schemes with respect to the linear arrangement of the average information contribution values of $E_{G}$:
@@
    Complete - all the edges in $I$ have grouped $G_{1}$ and $G_{2}$’s edges respectively.
@@
Averaging of information contribution values has been done because different vertex permutations result in different $K(G)$ values. Since $\Lambda(G)$ groups isomorphisms of different vertex permutations that are automorphic per grouping, $\Cup(G)$ is the compressed version of $\Lambda(G)$ and ultimately of $S_{G}$. Pairing this approach of getting unique permutations and averaging of each edge’s information contribution in a graph allows us to have a fair assessment of $BDM$ by probabilistic and combinatorial means. We computed the average information contribution values with respect to the parameter $\Gamma$ in $\mathfrak{C}(G,\Gamma)$ to determine if $\Cup(G)$ is already effective without using $S_{G}$ since $\Cup(G)$ is algorithmically quicker to execute. Tables 2 and 3 show a comparison of $\Cup(G)$ and $S_{G}$. We can see that the symmetric group is more effective compared to using automorphic subset sampling. Although using $\Cup(G)$ has partially grouped edges of some graphs, $S_{G}$ correctly grouped edges for more graphs.
@@
The connecting edge does not have $\max_{\text{info}}$ using this permutation set. Let $e_{n}$ be the connecting edge which is placed in the $n$th location in $I_{k}=\{e_{n+1},e_{n},e_{n-1},\ldots,e_{1}\}$, where $I_{k}$ is the set of keys for the set of tuples $I$ and $I[e_{n}]$ is the average information contribution value of the connecting edge. If the connecting edge is $\max_{\text{info}}$, then $n=|I|$ (the first value in the linear arrangement).
As seen in Table 4, not only did $\mathfrak{C}$ identify that the connecting edge was $\max_{\text{info}}$} for most graphs using both $\Gamma=\Cup(G)$ and $\Gamma=S_{G}$, but it also managed to identify the distance of the connecting edge to the next edge in the linear arrangement of average information contribution was more than the cutoff for half of the graphs. Out of the 30 graphs, 16 of them satisfied $e_{n}-e_{n-1}>\log_{2}(2)$. Interestingly, it shows that performing averaging of information contribution with either $\Cup(G)$ or $S_{G}$ was effective when graphs contain a sub-graph that has a high edge count. Among these 16 graphs, either a Watts-Strogatz or a Complete graph was present as a sub-graph for each graph. This indicates that $BDM$ is effective in detecting tightly connected substructures within a graph. Further proof is that when a graph only contains these two subgraphs, the connecting edge’s average information contribution is high relative to the other graphs examined.
Conversely, $BDM$ did not satisfy $e_{n}-e_{n-1}>\log_{2}(2)$ for 14 out of the 30 graphs (colored red in Table 4) when $\Gamma=\Cup(G)$. 6 of the 9 graphs have a random sub-graph that is included per graph. When both subgraphs are random graphs, then adding a connecting edge will not be distinguishable from the other edges in the graph because the entire graph would be considered one combined random graph. There was not much change in $e_{n}-e_{n-1}$ when using $S_{G}$ over $\Cup(G)$ except for graphs that had a cycle sub-graph. This indicates that averaging of information contribution using $S_{G}$ provides a more fine-grained inspection of substructures. Appending to the findings in the Edge Grouping section of the results, most of the graphs that were classified as complete or partial edge grouping consist of mostly graphs with regular and small-world subgraphs, namely Cycle, Complete and Watts-Strogatz graphs.
@@ Discussion
There are 2 limitations that prohibit us from further experimentation of larger graphs. First is that $\mathfrak{C}$ grows by $n!$, where $n$ is the number of nodes of the graph. To fully test $BDM$ on the basis of combinatorial and probabilistic means, much bigger graphs should be tested using $\mathfrak{C}$. Although $\mathfrak{C}$ shows promise of identifying substructures in a graph and was able to:
@@
it must be tested in real world networks whose number of nodes is greater than 12 to have a more accurate testing of $BDM$. The second limitation is in the sub-matrix size that $CTM$ (within $BDM$) can slice an adjacency matrix of a graph. Currently, $BDM$ can handle $3\times 3$ or $4\times 4$ sub matrices and uses the periodic partitioning when $n$ is not divisible by $l$. Ideally, since $BDM$ is an approximated upper-bound of algorithmic complexity, we would like for $CTM$ to handle $l>4$ for larger graphs since there is an error overhead whenever $n$ is not divisible by $l$ and thus the remaining rows and columns in the adjacency matrix are appended with the previous rows and columns using a partitioning technique. But this is a monumental task since computing for larger matrices in $CTM$ is equivalent to running the Busy Beaver function.
@@ Conclusion
A combination of permuting a graph then averaging the information contribution per edge perturbed was done to 30 synthetic graphs to determine the efficacy of $BDM$ as a method in revealing substructures within complex networks. The sets of permutations that the algorithm $\mathfrak{C}$ (which uses $BDM$) used was the automorphic subsets $\Cup(G)$ of a graph and its symmetric group $S_{G}$. We used these two permutation sets to compare a compressed representation of a graph’s structure (automorphic subsets) from its entire permutation set (symmetric group). Out of the 30 graphs, 29 of them have been successfully identified as having the connecting edge as the edge with the highest average information contribution. Within these 29 graphs, 16 were also identified of having the connecting edge as the edge with a distance of more than $\log_{2}(2)$ from other edges within their respective graphs. To highlight, edge perturbation was proven to be effective for graphs that have tightly connected subgraphs, namely Watts-Strogatz graphs and Complete graphs. The connecting edge was clearly differentiated in terms of its average information contribution from the edges of the subgraphs. Averaging of information contribution using $\Cup(G)$ and $S_{G}$ was also effective in completely or partially grouping edges together. Comparing the two permutation sets, $S_{G}$ performed better since more graphs’ edge grouping has classified as a complete edge grouping.
@@ Acknowledgments
I would like to acknowledge my thesis adviser Briane Samson for guiding and motivating me when writing this paper. I would also like to thank Hector Zenil for answering my inquiries regarding their initial implementation of causal deconvolution along with Allan Zea, who implemented the R package of causal deconvolution. Lastly, I would like to show my gratitude to the maintainer of the Python package that implements $BDM$, Szymon Talaga, for answering questions regarding implementing edge perturbation to the package.